# Relational Conformal Prediction for Correlated Time Series

### v1

@@ 3.2 Learning the Relational Structure
Assuming the dependency structure across time series to be unknown, we integrate a graph learning module into the architecture to derive the operational graph topology directly from the residuals. To do so, we adopt a probabilistic structure learning framework [39, 40, 41]. In particular, we associate each edge with a score $\phi^{ij}$ and learn a distribution over $K$-NN graphs parametrized by the matrix ${\bm{\Phi}}\in{\mathbb{R}}^{N\times N}$ [40, 42]. Notably, we consider graphs obtained by sampling, for each $i$-th node, $K$ elements without replacement from the categorical distribution
@@
where ${\mathcal{E}}_{\bm{\xi}}({}\cdot{})$ is a generic trainable encoder with parameters $\bm{\xi}$. In practice, sampling can be done efficiently by exploiting the GumbelTopK trick [43] and scores ${\bm{\Phi}}$ can be parametrized directly as ${\bm{\Phi}}=\bm{\xi}$.
@@
To propagate gradients through the sampling, we rely on the continuous relations introduced by  Xie and Ermon [44] paired with a straight-through gradient estimator [45] to obtain discrete samples. Optionally, we sparsify the gradients by backpropagating only through a random subset of the zero entries of ${\bm{A}}$ (more details are provided in App. D).
@@ 3.3 Theoretical analysis and further discussion
Let $P^{c}_{t+H}({\bm{X}}_{t+H})=p_{t+H}({\bm{X}}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ and $P^{c}_{\psi}({\bm{X}}_{t+H})=p_{\psi}({\bm{X}}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ be the true conditional data-generating distribution at the test point $t+H$ and the probability distribution associated with the learned quantile function ${\mathcal{Q}}_{\psi}$, respectively. Then
$P^{c}_{t+H}\left({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right)$
@@
The proof relies on the properties of the total variation of probability measures and can be found in App. A. Here, differently from the problem settings introduced in Sec. 2.1, we do not assume the process to be time-invariant. 3.1 links the conditional coverage gap to the approximation error in estimating the quantile function of the residuals. By making assumptions on the expressivity of the quantile regressor in Eq. 13 and on the stationarity process (e.g., by assuming a strongly mixing process), we can expect the total variation between the learned and true distribution to shrink asymptotically as the size of the calibration set increases. Moreover, in this case, monitoring the coverage gap on a validation set offers an estimate of the actual miscoverage on test data. Conversely, if we expect the process to be non-stationary, ${\mathcal{Q}}_{\psi}$ has to be updated over time to keep the coverage gap contained. Within this context, the next section discusses a simple and sample-efficient approach to make CoRel adaptive. Finally, the bound provided in 3.1 shares similarities with the one in [16], which bounds the miscoverage gap for CP from weighted empirical quantiles. 3.1 can be seen as an analogous result that holds when estimates obtained from empirical quantiles are replaced with a parametric function approximator.
3.4 Adaptativity
The RelQP model introduced in Sec. 3.1 can yield arbitrarily large coverage gaps in the presence of distribution shifts from the calibration set, where the model is trained. Adopting a re-training approach such as in Xu and Xie [19] would be impractical due to the higher sample complexity entailed by the deep learning approach that we adopt. Therefore, to mitigate this issue while keeping the computational complexity under control, we update only the local components of the model over time, i.e., the learnable node embeddings ${\bm{V}}$ [37]. This allows for keeping most of the parameters of the model fixed and fine-tuning only a small number of weights for each node. Empirically, we show that this procedure can effectively improve the quality of the uncertainty estimates.
@@ 4 Related Work
The problem of quantifying forecast uncertainty is central in fundamental and applied research in time series forecasting [9, 6]. Among deep learning approaches [1], many generative architectures have been proposed as means to obtain probabilistic forecasts [13, 46, 47, 48]. Most related to our approach are those methods that exploit quantile regression to produce probabilistic forecasts [27, 28, 49, 31]. Similarly to CoRel, these quantile regression techniques do not usually require strong assumptions on the underlying data distribution. Regarding probabilistic graph-based forecasting architecture, the existing literature is limited [4, 3]. Wu et al. [50] investigate the combination of STGNNs with standard uncertainty quantification techniques for deep learning. Pal et al. [51] use an STGNN to implement a state-space model and quantify uncertainty within a Bayesian framework. Wen et al. [52] propose a probability predictor based on combining STGNNs with a diffusion model [53]. Zambon et al. [54] introduce a framework for designing probabilistic graph state-space models that can process collections of time series. However, all these methods cannot operate on top of an existing pre-trained model and require training an ad-hoc forecasting model. Conversely, CoRel is trained, within a CP framework, on predicting the quantiles of the error distribution of an existing model, rather than on forecasting the target variable.
@@
Related work on CP for time series has been already discussed in Sec. 2.2 and Sec. 3. Related to our method, Mao et al. [55] propose a CP approach for (static) spatially correlated data. Jiang et al. [56] propose to quantify the uncertainty in predicting power outages by fitting a quantile random forest [25] on time series from neighboring geographical units. CoRel can be framed among the CP methods that learn a model of conformal scores distribution [19, 57]. Differently from existing methods that operate on each time series separately, the estimates are conditioned on errors at both the target time series as well as at neighboring nodes. To the best of our knowledge, no previous CP method has been designed to specifically operate on collections of correlated time series and exploit graph deep learning operators. CP methods for multivariate time series do exist [58, 59], but operate on a single multidimensional time series. Moreover, although global-local models are popular among forecasting architectures [36, 1], CoRel is the first CP architecture of this kind. Finally, CP methods have also been applied to static graphs and used to quantify the uncertainty of GNNs, both in inductive [60] and transductive settings [61]. More recently, Davis et al. [62] proposed a CP method for node classification with GNNs in dynamic networks. These methods often assume node/edge exchangeability [60, 61] or are limited to node classification tasks [63].
@@
Table 1: Performance comparison for $\alpha=0.1$. $\Delta$Cov values are color-coded to highlight different levels of undercoverage: green ($<$2%), yellow (2-3%), orange (3-4%), red ($>$4%). The lowest Winkler score for each scenario is shown in bold.
Metric SCP NexCP SeqCP SCPI HopCPT CoRNN CoRel
METR-LA RNN $\Delta$Cov -0.65 -1.07 -8.10 -0.58$\pm$0.01 -2.04$\pm$0.30 -1.71$\pm$0.60 -1.24$\pm$0.73
PI-Width 20.60 24.12 18.59 19.81$\pm$0.01 17.18$\pm$0.11 17.51$\pm$0.42 16.10$\pm$0.70
Winkler 40.22 39.19 48.29 37.27$\pm$0.01 28.41$\pm$0.20 33.90$\pm$0.21 25.55$\pm$0.44
Transf $\Delta$Cov -0.72 -1.22 -12.23 -0.59$\pm$0.01 -2.69$\pm$0.65 -1.24$\pm$0.39 -1.02$\pm$0.59
PI-Width 20.51 24.21 16.99 19.90$\pm$0.00 16.84$\pm$0.21 17.78$\pm$0.36 15.72$\pm$0.15
Winkler 40.29 39.23 49.07 37.66$\pm$0.01 28.58$\pm$0.36 33.99$\pm$0.21 25.27$\pm$0.15
STGNN $\Delta$Cov -1.61 -1.24 -11.84 -1.56$\pm$0.01 -2.31$\pm$0.29 -0.62$\pm$0.40 -1.54$\pm$1.01
PI-Width 17.68 22.20 13.68 16.59$\pm$0.01 16.07$\pm$0.11 15.94$\pm$0.35 15.40$\pm$0.52
Winkler 36.78 34.34 39.73 34.37$\pm$0.02 26.18$\pm$0.13 30.11$\pm$0.08 26.86$\pm$0.26
CER-E RNN $\Delta$Cov -3.42 0.12 -3.52 -3.41$\pm$0.01 -4.50$\pm$0.59 -1.31$\pm$0.44 -3.89$\pm$0.15
PI-Width 2.61 3.26 2.77 2.40$\pm$0.00 2.07$\pm$0.05 2.11$\pm$0.03 1.88$\pm$0.01
Winkler 5.71 5.46 5.80 5.31$\pm$0.00 4.12$\pm$0.06 4.13$\pm$0.03 3.82$\pm$0.03
Transf $\Delta$Cov -3.30 0.11 -3.60 -3.23$\pm$0.02 -4.51$\pm$0.53 -1.23$\pm$0.90 -3.34$\pm$0.34
PI-Width 2.55 3.18 2.70 2.37$\pm$0.00 2.04$\pm$0.07 2.08$\pm$0.04 1.90$\pm$0.02
Winkler 5.67 5.40 5.76 5.29$\pm$0.00 4.10$\pm$0.09 4.26$\pm$0.04 3.80$\pm$0.04
STGNN $\Delta$Cov -4.41 0.04 -3.63 -4.32$\pm$0.00 -5.28$\pm$0.17 -2.08$\pm$0.12 -5.03$\pm$0.52
PI-Width 2.29 2.98 2.42 2.10$\pm$0.00 1.84$\pm$0.01 1.90$\pm$0.03 1.80$\pm$0.04
Winkler 5.11 4.89 5.00 4.81$\pm$0.00 3.68$\pm$0.02 3.84$\pm$0.04 3.82$\pm$0.05
AQI RNN $\Delta$Cov 4.77 0.04 -2.34 3.80$\pm$0.02 3.97$\pm$1.07 0.47$\pm$1.60 -2.46$\pm$1.64
PI-Width 131.17 86.95 78.87 120.77$\pm$0.05 106.20$\pm$6.45 78.00$\pm$4.16 72.30$\pm$3.56
Winkler 159.23 132.59 136.64 152.17$\pm$0.05 133.87$\pm$2.39 111.21$\pm$2.11 109.79$\pm$1.53
Transf $\Delta$Cov 4.77 0.03 -2.23 3.85$\pm$0.01 4.21$\pm$0.43 -0.82$\pm$1.14 -1.04$\pm$1.07
PI-Width 131.91 87.45 80.15 122.81$\pm$0.03 108.38$\pm$2.70 75.10$\pm$1.67 74.44$\pm$1.47
Winkler 160.66 134.27 138.95 154.63$\pm$0.02 135.54$\pm$1.15 111.11$\pm$0.04 110.15$\pm$1.06
STGNN $\Delta$Cov 5.19 -0.01 -2.69 4.53$\pm$0.01 4.52$\pm$0.21 -0.74$\pm$0.61 -2.00$\pm$1.30
PI-Width 135.89 92.93 81.20 124.27$\pm$0.04 112.72$\pm$1.53 75.97$\pm$0.26 73.51$\pm$2.40
Winkler 163.12 139.80 141.37 154.26$\pm$0.04 138.96$\pm$0.58 112.90$\pm$0.29 112.62$\pm$0.91
@@ 5 Experiments
We validate CoRel across three experimental settings. In the first one (Sec. 5.1), we compare it against state-of-the-art CP methods operating on the residuals produced by different forecasting models. Then, we analyze CoRel in a controlled environment (synthetic dataset). Finally, we assess the effectiveness of the procedure described in Sec. 3.4 in adaptively improving the PIs. We implement CoRel as an RNN followed by two message-passing layers. To approximate the quantile function, we train the model by minimizing the pinball loss over a discrete set of quantiles, similarly to Wen et al. [27]. PIs are constructed as in Eq. 14; App. G shows results for the alternative construction in Eq. 16. To learn the graph, we directly parametrize the score matrix ${\bm{\Phi}}$ by associating a learnable parameter with each of its entries. We use as metrics: 1) the difference between the specified confidence level $1-\alpha$ and the observed coverage on the test set ($\Delta$Cov); 2) the width of the PI (PI-Width); 3) the Winkler score [64], which is computed as the width of the PI plus penalty for each observation outside of the predicted interval proportional to the actual error (Winkler). More details and results are provided in the appendix.

### v2

@@ 3.2 Learning the Relational Structure
Assuming the dependency structure across time series to be unknown, we integrate a graph learning module into the architecture to derive the operational graph topology directly from the residuals. To do so, we adopt a probabilistic structure learning framework (Niculae et al., 2023; Cini et al., 2023c; Manenti et al., 2024). In particular, we associate each edge with a score $\phi^{ij}$ and learn a distribution over $K$-NN graphs parametrized by the matrix ${\bm{\Phi}}\in{\mathbb{R}}^{N\times N}$ (Cini et al., 2023c; Kazi et al., 2022). Notably, we consider graphs obtained by sampling, for each $i$-th node, $K$ elements without replacement from the categorical distribution
@@
where ${\mathcal{E}}_{\bm{\xi}}({}\cdot{})$ is a generic trainable encoder with parameters $\bm{\xi}$. In practice, sampling can be done efficiently by exploiting the GumbelTopK trick (Kool et al., 2019) and scores ${\bm{\Phi}}$ can be parametrized directly as ${\bm{\Phi}}=\bm{\xi}$.
@@
To propagate gradients through the sampling, we rely on the continuous relations introduced by  Xie and Ermon (2019) paired with a straight-through gradient estimator (Bengio et al., 2013) to obtain discrete samples. Optionally, we sparsify the gradients by backpropagating only through a random subset of the zero entries of ${\bm{A}}$ (more details are provided in App. D). As already mentioned, different parametrizations and gradient estimators for subset samplers exist and can be considered. Furthermore, if the sparsity assumption is deemed unrealistic for the problem at hand, other distributions (e.g., based on Bernoulli random variables) can be considered (Cini et al., 2023c).
@@ 3.3 Theoretical Analysis and Further Discussion
Let $P^{c}_{t+H}({\bm{x}}^{i}_{t+H})=p_{t+H}({\bm{x}}^{i}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ and $P^{c}_{\psi}({\bm{x}}^{i}_{t+H})=p_{\psi}({\bm{x}}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ be the true conditional data-generating distribution at the test point $t+H$ and the probability distribution associated with the learned quantile function ${\mathcal{Q}}_{\psi}$, respectively. Then
$P^{c}_{t+H}\left({\bm{x}}^{i}_{t+H}\in\widehat{C}^{\alpha}_{i,t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right)$
@@
The proof relies on the properties of the total variation of probability measures and can be found in App. A. Here, differently from the problem settings introduced in Sec. 2.1, we do not assume the process to be time-invariant. 3.1 links the conditional coverage gap to the approximation error in estimating the quantile function of the residuals. The bound provided in 3.1 shares similarities with the one in (Barber et al., 2023), which bounds the miscoverage gap for CP from weighted empirical quantiles. 3.1 can be seen as an analogous result that holds when estimates obtained from empirical quantiles are replaced with a parametric function approximation. By making assumptions on the expressivity of the quantile regressor in Eq. 13 and on the stationarity process (e.g., by assuming a strongly mixing process), we can expect the total variation between the learned and true distribution to shrink asymptotically as the size of the calibration set increases. Moreover, in this case, monitoring the coverage gap on a validation set offers an estimate of the actual miscoverage on test data. Similar analyses have been carried out for recently introduced CP methods for time series (Xu and Xie, 2023b; Lee et al., 2025); we refer the reader to these related works. If we instead expect the process to be non-stationary, ${\mathcal{Q}}_{\psi}$ has to be updated over time to keep the coverage gap contained. Within this context, the next section discusses a simple and sample-efficient approach to make CoRel adaptive. Finally, the computational complexity of CoRel will depend on the STGNN used to implement the RelQP and the number of edges sampled while learning the relational structure. For example, the cost of a forward pass for a standard TTS STGNN with window size $W$ and a graph with $E$ edges would scale as $\mathcal{O}(WN+E)$ (Cini et al., 2023a).
3.4 Adaptation
The RelQP model introduced in Sec. 3.1 can yield arbitrarily large coverage gaps in the presence of distribution shifts from the calibration set where the model is trained. Adopting a re-training approach, such as in Xu and Xie (2023b), would be impractical due to the higher sample complexity entailed by the deep learning approach that we adopt. Therefore, to mitigate this issue while keeping the computational complexity under control, we update only the local components of the model over time, i.e., the learnable node embeddings ${\bm{V}}$ (Cini et al., 2023b). This allows for keeping most of the learnable parameters fixed and fine-tuning only a small number of weights for each node. Empirically, we show that this procedure can effectively improve the quality of the uncertainty estimates.
@@ 4 Related Work
The problem of quantifying forecast uncertainty is central in fundamental and applied research in time series forecasting (Hyndman and Athanasopoulos, 2018; Petropoulos et al., 2022). Among deep learning approaches (Benidis et al., 2022), many generative architectures have been proposed as means to obtain probabilistic forecasts (Salinas et al., 2020; Rangapuram et al., 2018; de Bézenac et al., 2020; Rasul et al., 2021). Most related to our approach are those methods that exploit quantile regression (Wen et al., 2017; Gasthaus et al., 2019; Kan et al., 2022; Gouttes et al., 2021). Similarly to CoRel, these quantile regression techniques do not usually require strong assumptions on the data distribution.
Uncertainty quantification in STGNNs
Regarding probabilistic graph-based forecasting architecture, the existing literature is limited (Jin et al., 2023; Cini et al., 2023a). Wu et al. (2021) investigate the combination of STGNNs with standard uncertainty quantification techniques for deep learning. Pal et al. (2021) use an STGNN to implement a state-space model and quantify uncertainty within a Bayesian framework. Wen et al. (2023) propose a probabilistic predictor based on combining STGNNs with a diffusion model (Ho et al., 2020). Zambon et al. (2023) introduce a framework for designing probabilistic graph state-space models that can process collections of time series. However, all these methods cannot operate on top of an existing pre-trained model and require training an ad-hoc forecasting model. Conversely, CoRel is trained, within a CP framework, on predicting the quantiles of the error distribution of any existing model, rather than on forecasting the target variable.
@@
Related work on CP for time series has been already discussed in Sec. 2.2 and Sec. 3. Related to our method, Mao et al. (2024) propose a CP approach for (static) spatially correlated data. Jiang et al. (2024) propose to quantify the uncertainty in predicting power outages by fitting a quantile random forest (Meinshausen and Ridgeway, 2006) on time series from neighboring geographical units. CoRel can be framed among the CP methods that learn a model of conformal scores distribution (Xu and Xie, 2023b; Lee et al., 2024). Unlike existing methods that operate on each time series separately, the estimates are conditioned on errors at both the target time series as well as at neighboring nodes. To the best of our knowledge, no previous CP method has been designed to specifically operate on collections of correlated time series and exploit graph deep learning operators. CP methods for multivariate time series do exist (Xu et al., 2024; Sun and Yu, 2024; Feldman et al., 2023), but operate on a single multidimensional time series. Moreover, although global-local models are popular among forecasting architectures (Smyl, 2020; Benidis et al., 2022), CoRel is the first CP architecture of this kind. Finally, CP methods have also been applied to static graphs and used to quantify the uncertainty of GNNs, both in inductive (Zargarbashi et al., 2023) and transductive (Huang et al., 2024) settings. These methods often assume node/edge exchangeability (Zargarbashi et al., 2023; Huang et al., 2024) or are limited to node classification (Clarkson, 2023) or link prediction (Zhao et al., 2024). Recently, Davis et al. (2024) proposed a CP method for node classification with GNNs in dynamic networks.
Metric SCP NexCP SeqCP SPCI HopCPT CoRNN CoRel
METR-LA RNN $\Delta$Cov -1.28 -1.00 -6.93 -1.24$\pm$0.01 -1.46$\pm$0.11 -0.49$\pm$0.53 -1.29$\pm$0.21
PI-Width 20.69 25.23 19.52 19.84$\pm$0.01 16.62$\pm$0.10 19.48$\pm$0.43 14.38$\pm$0.22
Winkler 40.80 41.33 50.12 37.95$\pm$0.01 25.63$\pm$0.19 30.24$\pm$0.18 23.78$\pm$0.20
Transf $\Delta$Cov -1.18 -0.97 -6.98 -1.13$\pm$0.00 -1.22$\pm$0.56 -0.55$\pm$0.61 -1.02$\pm$0.63
PI-Width 20.81 25.30 19.44 19.86$\pm$0.01 16.66$\pm$0.13 19.38$\pm$0.64 14.17$\pm$0.43
Winkler 40.55 41.64 49.90 37.74$\pm$0.01 25.37$\pm$0.33 30.33$\pm$0.18 23.65$\pm$0.19
STGNN $\Delta$Cov -0.99 -0.62 -13.60 -$0$0.13$\pm$0.27 -0.34$\pm$0.49 -0.92$\pm$0.31
PI-Width 17.30 22.20 12.87 16.38$\pm$0.01 15.75$\pm$0.19 16.20$\pm$0.24 14.65$\pm$0.24
Winkler 34.94 34.49 40.36 33.66$\pm$0.01 22.80$\pm$0.28 28.74$\pm$0.12 24.70$\pm$0.16
CER-E RNN $\Delta$Cov -3.46 0.11 -3.57 -3.45$\pm$0.00 -4.37$\pm$0.32 -2.24$\pm$0.34 -3.60$\pm$0.35
PI-Width 2.60 3.26 2.75 2.39$\pm$0.00 1.97$\pm$0.03 1.96$\pm$0.02 1.83$\pm$0.03
Winkler 5.69 5.48 5.79 5.29$\pm$0.00 3.87$\pm$0.07 3.84$\pm$0.01 3.71$\pm$0.04
Transf $\Delta$Cov -3.35 0.10 -3.53 -3.26$\pm$0.01 -3.95$\pm$0.60 -2.04$\pm$0.32 -3.97$\pm$0.24
PI-Width 2.52 3.16 2.67 2.33$\pm$0.00 2.01$\pm$0.09 1.94$\pm$0.03 1.80$\pm$0.02
Winkler 5.60 5.36 5.69 5.20$\pm$0.00 3.88$\pm$0.12 3.82$\pm$0.02 3.67$\pm$0.02
STGNN $\Delta$Cov -4.30 0.08 -3.83 -4.17$\pm$0.01 -5.06$\pm$0.15 -2.13$\pm$0.83 -4.99$\pm$0.61
PI-Width 2.28 3.00 2.42 2.09$\pm$0.00 1.79$\pm$0.01 1.85$\pm$0.04 1.77$\pm$0.05
Winkler 5.11 4.87 4.99 4.76$\pm$0.00 3.49$\pm$0.02 3.72$\pm$0.02 3.76$\pm$0.04
AQI RNN $\Delta$Cov 5.09 -0.63 -3.21 1.79$\pm$0.01 0.01$\pm$2.62 -2.06$\pm$1.52 -2.78$\pm$0.60
PI-Width 118.06 82.04 74.73 103.88$\pm$0.03 90.35$\pm$13.04 71.61$\pm$1.82 68.13$\pm$1.30
Winkler 148.61 131.18 135.59 143.10$\pm$0.01 133.24$\pm$8.31 113.11$\pm$0.77 107.67$\pm$0.94
Transf $\Delta$Cov 5.07 -0.66 -3.17 1.60$\pm$0.01 -2.34$\pm$1.19 -1.93$\pm$0.44 -1.81$\pm$1.67
PI-Width 118.50 81.15 75.08 104.11$\pm$0.01 80.37$\pm$2.09 73.25$\pm$0.95 72.25$\pm$3.16
Winkler 150.33 132.08 137.62 145.14$\pm$0.02 129.85$\pm$3.14 112.71$\pm$0.44 108.71$\pm$1.38
STGNN $\Delta$Cov 4.48 -0.32 -2.94 2.64$\pm$0.01 -0.78$\pm$1.34 -1.52$\pm$0.51 -2.00$\pm$1.54
PI-Width 111.68 80.01 71.92 99.90$\pm$0.03 79.39$\pm$3.38 70.17$\pm$0.80 68.24$\pm$2.65
Winkler 143.39 127.98 130.86 137.09$\pm$0.03 121.26$\pm$2.48 109.65$\pm$0.30 108.45$\pm$1.52
@@
Table 1: Performance comparison for $\alpha=0.1$. $\Delta$Cov values are color-coded: green (0-2%), yellow (2-3%), orange (3-4%), red ($>$4%). The lowest Winkler score for each scenario is shown in bold.
@@ 5 Experiments
We validate CoRel across three experimental settings. In the first one (Sec. 5.1), we compare it against state-of-the-art CP methods operating on the residuals produced by different forecasting models. Then, we analyze CoRel in a controlled environment (synthetic dataset). Finally, we assess the effectiveness of the procedure described in Sec. 3.4 in adaptively improving the PIs. We implement CoRel as an RNN followed by two message-passing layers. To approximate the quantile function, we train the model by minimizing the pinball loss over a discrete set of quantiles, similarly to Wen et al. (2017). PIs are constructed as in Eq. 14; Sec. G.2 shows results for the alternative construction in Eq. 16. To learn the graph, we directly parametrize the score matrix ${\bm{\Phi}}$ by associating a learnable parameter with each of its entries. We use as metrics: 1) the difference between the specified confidence level $1-\alpha$ and the observed coverage on the test set ($\Delta$Cov), 2) the width of the PI (PI-Width), and 3) the Winkler score (Winkler, 1972), which is computed as the width of the PI plus penalty for each observation outside of the predicted interval proportional to the actual error (Winkler). Note that balancing coverage and PI width is the main challenge. More details and results are provided in the appendix.